# HabiCrowd: A High Performance Simulator for Crowd-Aware Visual Navigation

### v1

@@ 3.2 Dataset Construction
4 Experiments
4.1 HabiCrowd Evaluation
@@
[…]
Figure 8: Performance and memory benchmarks.
Figure 9: Human poses comparison. Our normalization process ensures that all human poses have the upper body facing the front (a).
Human Dynamics Evaluation. We utilize the UMANS engine van2020generalized to compare our HabiCrowd dynamic model UPL++ with ORCA dynamic model in iGibson-SN on the following aspects: i) Average computational time (mCT): determines the average computational time that each dynamics model needs to update the state for each virtual human during the navigation process. ii) Collision avoidance rate (CAR): measures how many times collisions between virtual humans are avoided. iii) Goal-reaching rate (GR): computes the frequency of virtual humans navigating to their destinations.
Table 11 reports the evaluation results. This table shows that both iGibson-SN and our simulator achieve collision-free navigation. Although HabiCrowd demonstrates a goal-reaching rate comparable to iGibson-SN’s (with a slight margin of 0.78%), our model shows remarkable improvement in terms of computational time, which is nearly twice faster than iGibson-SN.
Recall from Section 3.1 that the computational complexity of iGibson-SN and HabiCrowd is $O(n)$. However, from Figure 11, we observe that the slope of HabiCrowd is considerably steeper than iGibson-SN’s, indicating that the corresponding constant factor of HabiCrowd is significantly smaller than its counterpart. As a result, the computational efficiency advantage of HabiCrowd over iGibson-SN is preserved even when the number of humans scales up.
Figure 10: Average computational time of human dynamics.
[…]
CAR $\uparrow$ GR $\uparrow$ mCT $\downarrow$ iGibson-SN 2021igibsonchallenge 100.00 94.09 0.0228 HabiCrowd (ours) 100.00 93.31 0.0117
Figure 10: Average computational time of human dynamics.
Figure 11: Human dynamics evaluation. We evaluate human dynamics models on the floor plans of HabiCrowd.
User Study. We provide a user study with 28 participants and ask them to rate HabiCrowd, Isaac Sim, and iGibson-SN on four criteria: crowdedness, human model realism, scene realism, and daily familiarity. Figure 7 shows that in all aspects, our simulator is preferred over iGibson-SN. Moreover, our simulator exhibits comparable metrics in terms of human model realism and scene realism compared to Isaac Sim. Our provided environments are also considered to be more crowded and daily familiar than Isaac Sim.
Human Poses Comparison. We extract human poses from all datasets by using the PCT approach geng2023human and provide the qualitative analysis in Figure 9. The findings reveal that our simulator showcases a greater degree of diversity in human poses when compared to the other simulators. Consequently, HabiCrowd captures a wider range of human pose variations, therefore, reflecting a more comprehensive representation of real-life human settings.
4.2 Crowd-aware Visual Navigation
@@
Experimental Settings. We examine two crowd-aware visual navigation tasks upon HabiCrowd, namely, point navigation wijmans2019dd and object navigation chaplot2020object . The agent is initialized at a pre-defined location in a human-crowded scene and asked to navigate to an objective. At each timestep, the agent receives RGB and depth images that depict its egocentric viewpoint. We introduce an additional objective that requires the agent to minimize the frequency of collisions with virtual humans. Each collision is counted when the skin-to-skin distance between the agent and any human zeros.
@@
Denote $d_{t}$ as the distance from the agent to the goal at timestep $t$, we design the reward model as: $r_{t}=R^{\text{c}}\mathbb{I}^{\text{c}}_{t}+R^{\text{s}}\mathbb{I}^{\text{s}}_{t}+R^{\text{f}}\times(d_{t}-d_{t-1})$, where $R^{\text{c}}$ is the collision penalty, $R^{\text{s}}$ is the reward for completing the task, $R^{\text{f}}$ is the shaping reward for moving closer to the objective, and $\mathbb{I}^{\text{c}}_{t}$, $\mathbb{I}^{\text{s}}_{t}$ are the indicators of collision and success of the current step, respectively. This reward model helps balance between navigating towards the objective and avoiding collisions. We train all methods using the above reward model. Followed by chaplot2020object , we set $R^{\text{s}}=2.5,R^{\text{f}}=1.0$. The default value of $R^{\text{c}}$ is set to $-10^{-4}$. For the human dynamics, we set $\tau^{\text{adj}}=0.5,\tau^{\text{soc}}=3.0,\tau^{\text{rot}}=0.2$ and $k^{\text{soc}}=1.5$ in all setups. Each method is trained in an end-to-end fashion using a total of 30M frames. We run each experiment on a computing cluster utilizing a node of four 48GB-RAM V100 GPUs.
@@
Baselines. We compare the following baselines: i) DS-RNN liu2021decentralized : A spatial and temporal learning-based agent. ii) DD-PPO wijmans2019dd : A widely-used state-of-the-art optimization algorithm in E-AI tasks. iii) LB-WPN tolani2021visual : A method that combines a learning-based perception module and a model-based planning module for autonomous navigation among humans. iv) DOA yokoyama2022benchmarking : The winning entry of the 2021 iGibson-SN challenge 2021igibsonchallenge that applies the dynamic obstacle augmentation method. v) OVRL-v2 yadav2023ovrl : A recent state-of-the-art baseline utilizing the vision transformer approach (ViT).
@@
Metrics. We employ three widely used metrics in traditional visual navigation tasks chaplot2020object , namely, i) Distance to goal (DTG), ii) Success rate (SR - in %), iii) Success path length anderson2018evaluation (SPL - in %). Additionally, we introduce iv) Collisions per distance (CPD), a new metric to quantify collisions occurring between the navigating agent and virtual humans. This metric is computed by dividing the number of collisions between the agent and humans by the distance traveled by the agent.
@@
Point Navigation Results. We present the point navigation results in Table 2. Overall, every baseline achieves decent performance, OVRL-v2 achieves the most comprehensive performance in terms of all metrics. Since HabiCrowd does not provide dynamics information regarding virtual humans to the agent due to our aim of enabling more realistic navigation settings, baselines such as DS-RNN liu2021decentralized , LB-WPN tolani2021visual are unable utilize human dynamics models to address the task effectively. On the other hand, OVRL-v2 leverages ViT, a robust vision backbone capable of extracting entities (such as virtual humans) from visual inputs as additional clues, thereby proficiently handling the task.
[…]
@@
Object Navigation Results. Table 3 illustrates the results on the HabiCrowd dataset by all approaches. It is obvious that object navigation poses a significantly greater challenge than point navigation, as evidenced by the decline in evaluation metrics for all baseline methods. The decrease in performance is aligned with the expectations outlined in khandelwal2022simple . The results show that OVRL-v2 has the best success rate, outperforming the second-best algorithm (DOA) by $3.523\%$.
@@
How does human density affect crowd navigation? Table 4 shows the success rates when we test all methods with different human density setups. OVRL-v2 stably outperforms other algorithms by $2.679\%$, $1.758\%$, and a significant margin of $2.924\%$ when the human density are from $<0.2$; $0.2-0.3$; and $>0.3$, respectively. In general, we observe that the higher the human density, the more challenging our task becomes as the success rates obtained by all methods decrease.
Table 4: Human density analysis. We show object navigation success rates of all methods by different human density setups.
Density $<0.2$ $0.2-0.3$ $>0.3$
LB-WPN tolani2021visual 7.363 3.328 1.515
DOA yokoyama2022benchmarking 8.416 4.337 2.201
DD-PPO wijmans2019dd 8.135 6.064 2.525
DS-RNN liu2021decentralized 7.854 5.051 2.530
OVRL-v2 yadav2023ovrl 11.095 7.822 5.454
[…]
Table 4: Human density analysis. We show object navigation success rates of all methods by different human density setups.
[…]
Table 5: Impact of collision penalty. We vary $R^{\text{c}}$ and showcase the object navigation performance.
@@
Does our reward model help? Table 5 reports the results of when all baselines use our reward model and when they do not. Evidently, our reward model yields improved performance in terms of CPD, SPL, and SR across all methods. We explain as follows. The introduced quantity $R^{\text{c}}$ contributes to agents’ regularization towards avoiding collisions with virtual humans, consequently improving performances in unseen cases.
@@
How does collision penalty affect crowd navigation? We study the impact of $R^{\text{c}}$ on our navigation problem. We observe in Table 5 that starting from $R^{\text{c}}=-10^{-4}$ to $-10^{-2}$, all baselines experience a downward tendency in success rates and an upward tendency in collision avoidance. This observation is to be expected, considering that $R^{\text{c}}$ serves as a regularizing factor for the collision avoidance component. Consequently, as $R^{\text{c}}$ diminishes, the focus of the baselines shifts towards prioritizing collision avoidance with virtual humans rather than navigating towards the goal.
Will attention mechanism help? Figure 12 demonstrates how the attention mechanism works in ViT backbone of OVRL-v2. We can see that the agent trained with the transformer-based method learns to attend to virtual humans of the scene to extract more useful information to complete the navigation tasks. However, it is noticeable that ViT also allocates attention to image patches representing the wall and ceiling, indicating potential areas for improvements of the attention mechanism.
@@
What is the drawback of current methods? A notable weakness of all current baselines is that they do not possess an explicit strategy to determine human dynamics from RGB-D egocentric observations. Although OVRL-v2 has a robust vision backbone to extract human dynamics information, the employed mechanism remains implicit and demands for improvements.
@@
Discussion. From the experiments, we can see that the point-goal crowd-aware navigation task is well-addressed, while the object-goal task remains a challenge. We have highlighted the critical issue in 3D crowd-aware visual navigation tasks, emphasizing the necessity of explicitly extracting human dynamics solely from RGB-D observations. This direction should be studied as making efforts towards solving crowd-aware navigation in real-life settings, where robots can only perceive human behaviors via sensor inputs without the assumption of a known human dynamics model.
5 Conclusion and Future Work
@@
We have presented HabiCrowd, a new crowd-aware simulator built upon Habitat 2.0. HabiCrowd has a robust human dynamic model and a diverse range of virtual humans compared to related simulators. The intensive experimental results indicate that our simulator utilizes computational resources more efficiently than its counterparts. We evaluate two visual navigation tasks and show that the recognition of virtual humans depicted in egocentric inputs is essential. Nevertheless, HabiCrowd still has limitations. While humans exhibit natural movements in daily lives, the constraints imposed by Habitat 2.0 szot2021habitat only allow for rigid movements. This assumption falls considerably behind the complexity of real-life practice and thus requires further enhancements. Finally, applying the learned policy from Habicrowd to the real world remains an open problem for future work.

### v2

@@ III-B Dataset Construction
Figure 7: Computational time.
IV Experiments
IV-A HabiCrowd Evaluation
@@
Figure 9: Human poses comparison.
Human Dynamics Evaluation. We leverage UMANS engine [42] to compare our dynamic model UPL++ with ORCA dynamic model in iGibson-SN on the following metrics: i) Mean computational time (mCT), the average time each model takes to update states for virtual humans during navigation. ii) Collision avoidance rate (CAR), measuring how often collisions are prevented between virtual humans. iii) Goal-reaching rate (GR), indicating the frequency of virtual humans reaching their destinations Table IV reports the evaluation results. This table shows that both iGibson-SN and our simulator achieve collision-free navigation. Although HabiCrowd demonstrates a goal-reaching rate comparable to iGibson-SN’s (with a slight margin of 0.78%), our model shows remarkable improvement in terms of computational time, which is nearly twice faster than iGibson-SN.
Recall from Section III-A that the computational complexity of iGibson-SN and HabiCrowd is $O(n)$. However, from Figure 7, we observe that the slope of iGibson-SN is considerably steeper than HabiCrowd’s, indicating that the corresponding constant factor of HabiCrowd is significantly smaller than its counterpart. As a result, the computational efficiency advantage of HabiCrowd over iGibson-SN is preserved even when the number of humans scales up.
TABLE III: Human Dynamics Evaluation.
CAR $\uparrow$ GR $\uparrow$ mCT $\downarrow$
iGibson-SN [17] 100% 94.1% 0.023
HabiCrowd (our) 100% 93.3% 0.012
TABLE IV: Human Density Analysis.
Baseline Density <0.2 0.2-0.3 >0.3
LB-WPN [43] 16.37 12.64 6.28
DOA [19] 17.78 13.25 7.36
DD-PPO [4] 15.83 11.79 7.07
DS-RNN [9] 14.95 7.51 4.24
OVRL-v2 [1] 22.38 16.49 11.27
User Study. We provide a user study with 52 participants from diverse professions, such as designers, animators, developers, and marketers. We ask them to rate HabiCrowd, Isaac Sim, and iGibson-SN on four criteria: crowdedness, human model realism, scene realism, and daily familiarity. Figure 7 shows that in all aspects, our simulator is preferred over iGibson-SN. In addition, HabiCrowd shows comparable levels of human model realism and scene realism, with a higher familiarity compared to Isaac Sim.
Human Poses Comparison. We extract human poses from all datasets by using the PCT approach [44] and provide the qualitative analysis in Figure 9. The findings reveal that our simulator showcases a greater degree of diversity in human poses when compared to the other simulators. Consequently, HabiCrowd captures a wider range of human pose variations, therefore, reflecting a more comprehensive representation of real-life human settings.
TABLE V: Visual Navigation Results.
Point Navigation Object Navigation
CPD $\downarrow$ SPL $\uparrow$ SR $\uparrow$ DTG $\downarrow$ CPD $\downarrow$ SPL $\uparrow$ SR $\uparrow$ DTG $\downarrow$
LB-WPN [43] 0.0521 78.13 91.84 0.6048 0.0462 7.989 15.31 5.801
DOA [19] 0.0626 80.44 92.96 0.5809 0.0441 8.278 16.57 5.742
DD-PPO [4] 0.0617 79.34 92.55 0.5991 0.0415 7.424 14.93 5.956
DS-RNN [9] 0.0476 76.82 90.73 0.6674 0.0305 6.078 12.34 5.972
OVRL-v2 [1] 0.0322 82.65 94.61 0.4895 0.0229 10.267 20.79 5.413
TABLE VI: Impact of Collision Penalty.
Without our reward model With our reward model
$R^{\text{c}}=0.0$ $R^{\text{c}}=-10^{-4}$ $R^{\text{c}}=-10^{-3}$ $R^{\text{c}}=-10^{-2}$
@@
LB-WPN [43] 0.076 5.90 10.62 6.19 0.046 7.99 15.31 5.80 0.036 4.78 8.99 6.38 0.021 3.44 6.04 6.78
DOA [19] 0.085 6.89 12.68 5.95 0.044 8.28 16.57 5.74 0.028 5.22 9.35 6.36 0.031 3.74 6.29 6.80
DD-PPO [4] 0.059 6.23 11.74 6.03 0.042 7.42 14.93 5.96 0.032 4.54 8.26 6.41 0.027 3.13 5.83 6.89
DS-RNN [9] 0.055 5.28 9.26 6.14 0.031 6.08 12.34 5.97 0.026 3.48 7.42 6.67 0.018 2.15 4.73 6.94
OVRL-v2 [1] 0.041 7.57 14.09 5.86 0.023 10.27 20.79 5.41 0.020 6.15 11.73 6.27 0.014 4.22 7.79 6.68
IV-B Crowd-aware Visual Navigation
@@
Experimental Settings. We examine two crowd-aware visual navigation tasks upon HabiCrowd, namely, point navigation [4] and object navigation [5]. The agent is initialized at a pre-defined location in a human-crowded scene and asked to navigate to an objective. At each timestep, the agent receives RGB and Depth observations. We introduce a new objective that requires the agent to minimize the number of collisions with virtual humans.
@@
Denote $d_{t}$ as the distance from the agent to the goal at timestep $t$, we design the reward model as: $r_{t}=R^{\text{c}}\mathbb{I}^{\text{c}}_{t}+R^{\text{s}}\mathbb{I}^{\text{s}}_{t}+R^{\text{f}}\times(d_{t}-d_{t-1})$, where $R^{\text{c}}$ is the collision penalty, $R^{\text{s}}$ is the reward for completing the task, $R^{\text{f}}$ is the shaping reward for moving closer to the objective, and $\mathbb{I}^{\text{c}}_{t}$, $\mathbb{I}^{\text{s}}_{t}$ are the indicators of collision and success of the current step, respectively. This reward model helps balance between navigating towards the objective and avoiding collisions. We train all methods using the above reward model. Followed by [5], we set $R^{\text{s}}=2.5,R^{\text{f}}=1.0$. The default value of $R^{\text{c}}$ is set to $-10^{-4}$. For the human dynamics, we set $\tau^{\text{adj}}=0.5,\tau^{\text{soc}}=3.0,\tau^{\text{rot}}=0.2$ and $k^{\text{soc}}=1.5$ in all setups. Each method is trained in an end-to-end fashion using a total of 30M frames. We run each experiment on a computing cluster utilizing a node of four 48GB-RAM V100 GPUs.
@@
Baselines. We compare the following baselines: i) DS-RNN [9]: A spatial and temporal learning-based agent. ii) DD-PPO [4]: A widely-used state-of-the-art optimization algorithm in E-AI tasks. iii) LB-WPN [43]: A method that combines a learning-based perception module and a model-based planning module for autonomous navigation among humans. iv) DOA [19]: The winning entry of the 2021 iGibson-SN challenge [17] that applies the dynamic obstacle augmentation method. v) OVRL-v2 [1]: A recent baseline utilizing the vision transformer approach (ViT).
@@
Metrics. We employ three metrics in traditional visual navigation tasks [5], namely, i) Distance to goal (DTG), ii) Success rate (SR - in %), iii) Success path length [45] (SPL - in %). Additionally, we introduce iv) Collisions per distance (CPD), a new metric to quantify collisions occurring between the navigating agent and virtual humans. This metric is computed by dividing the number of collisions between the agent and humans by the distance traveled by the agent.
@@
Point Navigation Results. We present the point navigation results in Table V. Overall, every baseline achieves decent performance, OVRL-v2 achieves the most comprehensive performance in terms of all metrics. Since HabiCrowd does not provide dynamics information regarding virtual humans to the agent due to our aim of enabling more realistic navigation settings, baselines such as DS-RNN [9], LB-WPN [43] are unable utilize human dynamics models to address the task effectively. On the other hand, OVRL-v2 leverages ViT, a robust vision backbone capable of extracting entities (such as virtual humans) from visual inputs as additional clues, thereby proficiently handling the task.
@@
Object Navigation Results. Table V illustrates the results on the HabiCrowd dataset by all approaches. It is observable that object navigation poses a significantly greater challenge than point navigation, as evidenced by the decline in evaluation metrics for all baseline methods. The decrease in performance is aligned with the expectations outlined in [25]. The results show that OVRL-v2 has the best success rate, outperforming the second-best algorithm (DOA) by $4.22\%$.
@@
How does human density affect crowd navigation? Table IV shows the success rates when we test all methods with different human density setups. OVRL-v2 stably outperforms other algorithms by $4.60\%$, $3.24\%$, and $3.91\%$ when the human density are from $<0.2$; $0.2-0.3$; and $>0.3$, respectively. In general, we observe that the higher the human density, the more challenging our task becomes as the success rates of all methods decrease.
@@
Does our reward model help? Table VI reports the results with and without our reward model for all baselines. Clearly, our reward model improves CPD, SPL, and SR performance metrics across all methods. We explain as follows. The introduced quantity $R^{\text{c}}$ contributes to agents’ regularization towards avoiding collisions with virtual humans, consequently improving performances in unseen cases.
@@
How does collision penalty affect crowd navigation? We study the impact of $R^{\text{c}}$ on our navigation problem. We observe in Table VI that starting from $R^{\text{c}}=-10^{-4}$ to $-10^{-2}$, all baselines experience a downward tendency in success rates and an upward tendency in collision avoidance. This observation is to be expected, considering that $R^{\text{c}}$ serves as a regularizing factor for the collision avoidance component. Consequently, as $R^{\text{c}}$ diminishes, the focus of the baselines shifts towards prioritizing collision avoidance with virtual humans rather than navigating towards the goal.
@@
What is the drawback of current methods? A notable weakness of all current baselines is that they do not possess an explicit strategy to determine human dynamics from RGB-D egocentric observations. Although OVRL-v2 leverages a robust vision backbone to extract human dynamics, the employed mechanism remains implicit and requires refinement.
V Discussion
@@
From the experiments, we can see that the point-goal crowd-aware navigation task is well-addressed, while the object-goal task remains a challenge. We have highlighted the critical issue in 3D crowd-aware visual navigation tasks, emphasizing the necessity of explicitly extracting human dynamics solely from RGB-D observations. This direction should be studied as making efforts towards solving crowd-aware navigation in real-life settings, where robots can only perceive human behaviors via sensor inputs without the assumption of a known human dynamics model.
Training navigation agents in simulations like Habitat shows promise, yet transferring these learned policies to real-world environments, such as on robotic platforms, continues to be a significant challenge [46]. Early-stage sim2real initiatives, including ROS interfaces, confront issues like physics discrepancies [47]. Furthermore, the gap between simulated and practical environments is widened by differences in visual appearance, absence of real-world noise, and oversimplified physical interactions, leading to poor real-world generalization of simulation-learned navigation policies [48]. In line with these efforts, our work aims to mitigate these issues by incorporating human dynamics into the Habitat simulator to create a more realistic bridge between the simulated world and the practical environment. We believe integrating human dynamics into HabiCrowd represents an encouraging step towards developing robots capable of learning in dynamic, human-aware environments that more closely replicate the natural environment encountered daily.
VI Conclusion and Future Work
@@
We have presented HabiCrowd, a new crowd-aware simulator built upon Habitat 2.0. HabiCrowd has a robust human dynamic model and a diverse range of virtual humans compared to related simulators. The intensive experimental results indicate that our simulator utilizes computational resources more efficiently than its counterparts. We evaluate two visual navigation tasks and show that the recognition of virtual humans depicted in egocentric inputs is essential. Nevertheless, HabiCrowd still has limitations. While humans exhibit natural movements in daily lives, the constraints imposed by Habitat 2.0 [35] only allow for rigid movements. This assumption falls considerably behind the complexity of real-life practice and thus requires further enhancements.